OpenStudy (anonymous):

The third term of an arithmetic sequence is 9 and the seventh term is 31. Find the sum of the first twenty-two terms.

OpenStudy (qqstory):

what is al?

OpenStudy (qqstory):

al The third term?

OpenStudy (cwtan):

a+2d=9 a+6d=31 solve a and d. Sum of first nth term= (n/2)(2a+(n-1)d) sub the value inside and u will found the answer

OpenStudy (anonymous):

thats not supossed to be there

OpenStudy (amistre64):

7-3 = 4 the difference between the 7th and 3rd terms is: 31-9 = 22 22/4 = the common difference, if i did that right :)

OpenStudy (qqstory):

22/4= 5.5. wrong, the third term is 9.

OpenStudy (cwtan):

1st term is -2?

OpenStudy (amistre64):

9 + 4(22/4) = 31, well, is does pan out 3: 9 4: 36/4 + 22/4 = 58/4 5: 58/4 + 22/4 = 80/4 6: 80/4 + 22/4 = 102/4 7: 102/4 + 22/4 = 124/4 = 31

OpenStudy (amistre64):

an = a1 + d(n-1) a3 = a1 + 22/4 (3-1) 9 - 22/2 = a1 = -2 a31 = -2 + 22/4 (30) the sum of the terms is then: \[\frac{31(-2-2+\frac{22*30}{4})}{2}\]

OpenStudy (amistre64):

why did i read 22 as 31?

OpenStudy (amistre64):

simple enough change :) \[\frac{22(-2-2+\frac{22*21}{4})}{2}\]

OpenStudy (cwtan):

haha

OpenStudy (amistre64):

i learnt something new today that im bound to forget by tomorrow; its something about arithmetic means between 2 numbers ....